Five years ago A age was four times the age of B. Five years hence A age will be the age of B find their present ages.
Answer by Boreal(15235)

As the problem is stated there is no solution.
Five years ago, A's age was 4 times that of B. That makes him older than B.
Yet, in 5 years A will be the same age that B is now. That makes him younger than B. He can't be both.
Maybe 5 years ago A's age, *4, was equal to B
4(A-5)=B
in 5 years, A's age will equal B's.
A+5=B
then 4A-20=A+5, substituting.
3A=25
A=8 1/3
B=13 1/3
5 years ago, A was 3 1/3, or 13/3. That is 1/4 th of B's age now.